OpenStudy (anonymous):

IF \[4\sin+4\cos=0\] THEN FIND OTHER TRIGONOMETRIC RATIOS.

Directrix (directrix):

4 sin x + 4 cos x = 0 sin x + cos x = 0 sin x = -cos x The values of x which makes this true are additive inverses. Recall that sine and cosine are equal when x = 45 degrees. For these to be additive inverses, the angle x will lie in quadrants where x and y have different signs.

OpenStudy (anonymous):

but what will be other ratios

OpenStudy (anonymous):

cosec x sec x tan x cot x

OpenStudy (sirm3d):

first, what is \(x\) such that \[sin x = -\cos x\] once you have that value(s), evaluate the other functions \(\sin x, \cos x, \tan x, \cot x\, \sec x,\) and \(\csc x\)

OpenStudy (anonymous):

but tan tan x is coming cos/sin which should be sin/cos